two divided by the difference of a number and 3 minus 4 divided by a number plus 3, equals 8 times the reciprocal of the difference of the number squared and 9. what is the number?
Answer by lwsshak3(11628)

two divided by the difference of a number and 3 minus 4 divided by a number plus 3, equals 8 times the reciprocal of the difference of the number squared and 9. what is the number?
**
let x=number
2/(x-3)-4/(x+3)=8/(x^2-9)=8/(x+3)(x-3)
LCD: (x+3)(x-3)
2(x+3)-4(x-3)=8
2x+6-4x+12=8
-2x+18=8
-2x=-10
x=5